Grade 12Algebra

There are 20 girls and 15 boys in a class
What is the ratio of a number of girls to the total number of students in the class?

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

Given
Number of girls = 20 girls
Number of boys = 15 boys
Total number of students = 20 + 15
= 35
Ratio of number of girls to total number of students = 20 / 35 = 4 / 7

Thanks